OpenStudy (anonymous):

Simplify 4 times square root 24 minus 4 square root 8 plus square root of 98

OpenStudy (anonymous):

Here's a hint: \[4 \sqrt{24} - 4 \sqrt{8} + \sqrt{98} = 4 \sqrt{\color{red} 4 \cdot 6} + 4 \sqrt{\color{red} 4 \cdot 2} + 4\sqrt{\color{red}{49}\cdot 2}\]

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

@yakeyglee what is the answer? i dont get it